OpenStudy (anonymous):

How do you solve: the square root of 864x^2?

OpenStudy (anonymous):

u square root 864 on your calculator, and as for X^2 it goes to X

OpenStudy (anonymous):

Please show step by step.

OpenStudy (anonymous):

k well u cant really show step by step its just plugging into your calculator or using guess and check

OpenStudy (anonymous):

and express answer in lowest radical form with no decimal

OpenStudy (anonymous):

the answer is \[12x \sqrt{6}\]

OpenStudy (anonymous):

\[\sqrt{9*16*6*x^2}=3*4*x \sqrt{6}=12x \sqrt{6}\]

OpenStudy (anonymous):

probably should know that \[864=2^5\times 3^3\] then \[\sqrt{2^5\times 3^3\times x^2}=2^2\times 3\times x\times \sqrt{2\times 3}\]

OpenStudy (anonymous):

math! ur online! wanna chat? im so bored.

OpenStudy (anonymous):

well i got 12xsqrt6. cuz sqrt x^2 is x, and u can pull out a 16 and a 9 from 864. sqrt of 16 is 4, sqrt9 is 3.ur left with 6 in the sqrt. so 12xsqrt6